Billings Metals Corp. just paid a $4.40 dividend to its common stockholders. You expect these dividends to grow at a constant annual rate of 5% indefinitely. Assume the risk-free rate is 3% and the market risk premium is 7%. If Billings Metals has a beta of 0.8, what is the intrinsic value of its common stock?

Select one: a. $128.33 b. $46.20 c. none of the above d. $82.50
Answer #1
Intrinsic Value = a. $128.33  
Dear Student
Thank you for using Chegg
Please find below the answer
Statementshowing Computations
Paticulars
ke = Rf + Beta (MRP)
ke = 3% + .80*7%
ke = 3% + 5.6%
ke = 8.6%
Intrinsic value P0 = D1/(ke-g)
P0 = 4.4*1.05/(8.6%-5%)
P0 = 4.62/(3.6%)
P0 = $128.33